OpenStudy (anonymous):

HELP! Which set of integers is a Pythagorean triple? A. 2, 3, 6 B. 3, 4, 5 C. 4, 5, 9 D. 5, 7, 12

OpenStudy (campbell_st):

well you can test them using the theorem is the longest side squared equal to the sum of the shorter sides squared (a) is \[6^2 = 3^2 + 2^2\] 36 = 9 + 4....? if its true that's the answer... now try it with the other 3 choices.